If f(x)= 8/(x^(1/2)) - 6x^(-2) - 8x
What is the derivative of f(x)?

Answer by stanbon(75887)

If f(x)= 8/(x^(1/2)) - 6x^(-2) - 8x
------------------------
f(x) = 8(x^(-1/2)) - 6x^(-2) - 8x
----
f'(x) = (-4)(x^-(3/2)) + 12x^(-3) - 8
----
f'(x) = -4/x^(3/2) + 12/x^3 - 8
=================================
Cheers,
Stan H.
==========